The springs of a 1700 kg car compress 4.0 mm when its 64 kg driver gets into the driver's seat.

If the car goes over a bump, what will be the frequency of oscillations? Ignore damping.

Express your answer to two significant figures and include the appropriate units.
Answer #1

First, find the spring constant of the springs

Given that the springs of the car compress 4mm when its 64 kg driver gets into the driver's seat.

Use Hooke's Law

F=kx

W=kx

mg=kx

\frac{mg}{x}=k

\frac{64kg*9.81m/s^{2}}{0.004m}=k

\frac{64*9.81}{0.004}=k

{\color{Red} k=156960N/m}

====================

Frequency, f=\frac{1}{2\pi }\sqrt{\frac{k}{m}}

f=\frac{1}{2\pi }\sqrt{\frac{k}{m_{car}+m_{driver}}}

f=\frac{1}{2\pi }\sqrt{\frac{156960N/m}{1700kg+64kg}}

f=\frac{1}{2\pi }\sqrt{\frac{156960}{1764}}

ANSWER: {\color{Red} f=1.5013Hz}

==========================